Shift report 20020112 00:00-08:00

From: Apache (apache@pii3.brahms.bnl.gov)
Date: Sat Jan 12 2002 - 07:59:25 EST

  • Next message: Apache: "Shift report 20020114 08:00-16:00"

    ------------------------------------------------------------
    Date: Sat Jan 12 07:59:25 2002, Shift: 00:00-08:00
    Supervisor: mads, with: (IGB)
    Log book#12 pages 49 - 50
    
    Shift Summary:
    ------------------------------------------------------------
    Had a long continuous store wich lasted
    from 2330 to 0515, at wich point the beam
    was suddenly lost due to Power Dip.
    Took dataruns 6315 through 6319 during this 
    period and later pedestal runs 6421-6423
    
    At 0000 there occured a problem with the  
    ITD server, wich disabled brahvo to open the
    detector windows. It also disabled ssh to 
    any server at bnl. During this period, data 
    was taken continously in one long run. (6418)
    As instructed by JH.
    
    The polarities on D1-D4 were changed from A to 
    B during the night, and are currently at
    1/3 B. 
    
    There was also problems with the printer, and 
    i could not print supermon outputs. (could 
    someone fix this?)
    ------------------------------------------------------------
    



    This archive was generated by hypermail 2b30 : Mon Jan 14 2002 - 11:32:13 EST